OpenStudy (samigupta8):

The number of decreasing function f:R-->R such that f(f(x))=x+1 for all x€R

Parth (parthkohli):

Let \(f(x)\) be \(ax+b\) where \(a<0\). Then \(f(f(x)) = a(ax+b)+b = a^2 x+ab+b \). The only possible value for \(a\) then has to be \(-1\) by comparing coefficients. But then there'd be no way to get the constant term as +1 'cuz \(ab+b = -1\cdot b + b = 0\).

OpenStudy (bobo-i-bo):

@ParthKohli That only says that f(x) isn't in the form ax+b. Doesn't mean there isn't another function which satisfies the conditions. I have a hypothesis which I don't know whether it's true, but I'm guessing it is, and it might help with the Q: If f:R->R is a function such that f(f(x)) is continuous, then f(x) is continuous.

OpenStudy (bobo-i-bo):

Maybe the condition that f(f(x)) is not only continuous but also bijective is necessary for my hypothesis.

OpenStudy (kainui):

Ahh ok I think I figured it out, the question specifically states decreasing, and nothing about continuous functions. So check this out: \[f(f(f(x))) = f(x)+1 = f(x+1)\] So since \(f(x+1)-f(x) >0\) all choice of f(x) are increasing. So none satisfy it.

OpenStudy (bobo-i-bo):

Nice!

OpenStudy (bobo-i-bo):

I'd still like to find out whether my hypothesis is true though :P

OpenStudy (kainui):

I didn't come across this so nicely as this, I had taken the derivative, looking for some kinda thing to do with it decreasing: \(f'(f(x))f'(x) = 1\) and recognized if I plugged in f(x) I'd get \(f'(x+1)f'(f(x))=1\) and have \(f'(x)=f'(x+1)\) and that's when I backpedaled haha... Alright let's see I'm trying to see if I can find a counterexample to your hypothesis first before I try to prove it haha >:D

OpenStudy (samigupta8):

Well did! Kainui sir .. Thanks.

Parth (parthkohli):

@Bobo-i-bo That's exactly what it means. Do give an example of a function which isn't of that form and still satisfies \(f(f(x))=x+1\)

OpenStudy (bobo-i-bo):

@ParthKohli Can't think of one! I don't know if one exists. :P But your argument does not prove that there does not exist a function which isn't of that form and still satisfies f(f(x))=x+1!

OpenStudy (kainui):

I had found a counterexample to the nonbijective case, but I think you sorta had that idea in mind and my computer died right before I hit post. I think my example was \(f(x) = \lfloor \sin(\pi x) \rfloor\) but yeah, clearly not bijective so I'm still wondering about this too. I feel like the definition of continuous maps as I learned it in real analysis has the answer in it something about composition of continuous maps etc etc but I'm too scared to look right now haha.